Polarised Light

Andrew Mark Allen - 05370299 April 8, 2012

Abstract

The proj TODO: this

Aims

TODO: this

Introduction and Theory

TODO: this

Experimental Method

TODO: this

Results and Analysis

Malus’ Law

According to Malus’ Law, the intensity of the transmitted light should vary with cos^{2}(θ)

The data

The experimental results for the Malus’ Law part of the experiment are shown in Figure 1. It can be seen how the

TODO: this

Discussion and Conclusion

TODO: This Figure 1: Plot of y^{2} against ring number

2